OpenStudy (anonymous):

Of the runners in a marathon through Northeastern Pennsylvania, 94.12% finished the race. If 2,125 runners started the race, how many finished the race?

Nnesha (nnesha):

94.12 divide by 100 first and then multiply by 2125

OpenStudy (aum):

\(\Large 2125 \times \frac{94.12}{100} = ?\)

OpenStudy (anonymous):

2,000 people finished the race

OpenStudy (anonymous):

To find a percentage of a number, multiply the number by the decimal version of the percent. For 94.12%, the decimal form is 0.9412 (because 94.12% really means 94.12 out of 100). So, to find 94.12% of 2125, multiply the decimal form by the number: 2125 × 0.9412 =2000.05, so 2000 people finished the race.

Nnesha (nnesha):

make ur post

Nnesha (nnesha):

do you see the *ask a question* box post your question there